OpenStudy (anonymous):

WIthout using a calculator, evaluate (lg 5)^2 + lg 2 lg 50

OpenStudy (raden):

(log 5)^2 + log 2 log 50 = (log 5)^2 + lg 2 (log 5 + log 10) = (log 5)^2 + lg 2 (log 5 + 1) = (log 5)^2 + lg 2 * log 5 + log 2 = (log 5)(log 5 + log 2) + log 2 = log 5 log 10 + log 2 = log 5 (1) + log 2 = log 5 + log 2 = log 10 = 1

OpenStudy (dumbcow):

that is correct, assuming you are using log base 10